OpenStudy (anonymous):

Solve for the variable: -15r – 8 = 37 A. -5 B. -3 C. -2 D. 3

OpenStudy (anonymous):

-15r-8+ 37 -15r=45 -r=3 r=3 D) 3

OpenStudy (anonymous):

i really appreciate your help thanks

OpenStudy (anonymous):

I disagree, I think it would be B: -15r-8=37; add 8 to both sides -15r=45; divide both sides by -15 in order to get the variable (x) by itself r=-3; That is the answer I think is right......

OpenStudy (anonymous):

my bad, left that negative out thnx